# Defensive Model Expansion for Robust Bayesian Inference

### v1

@@ Abstract
Some applied researchers hesitate to use nonparametric methods, worrying that they will lose power in small samples or overfit the data when simpler models are sufficient. We argue that at least some of these concerns are unfounded when nonparametric models are strongly shrunk towards parametric submodels. We consider expanding a parametric model with a nonparametric component $r(x)$ that is heavily shrunk toward zero. This construction allows the model to adapt automatically: if the parametric model is correct, the nonparametric component disappears, recovering parametric efficiency, while if it is misspecified, the flexible component activates to capture the missing signal. We show that this adaptive behavior follows from simple and general conditions. Specifically, we prove that Bayesian nonparametric models anchored to linear regression, including variants of Gaussian processes regression and Bayesian additive regression trees, consistently identify the correct parametric submodel when it holds and give asymptotically efficient inference for regression coefficients. In simulations, we find that the general BART model performs identically to correctly specified linear regression when the parametric model holds, and substantially outperform it when nonlinear effects are present. This suggests a practical paradigm: defensive model expansion as a safeguard against model misspecification.
@@ 1 Introduction
  • If the parametric model is badly misspecified or omits important signals, expanding into a nonparametric model can improve power. With a flexible “alternative” model where $r(x)\neq 0$, we can also use the estimated size of $r(\cdot)$ as a diagnostic for parametric model misspecification: if $|r(x)|$ is large on average then this is evidence that the parametric submodel model fits poorly.
@@ 1.1 Existing Literature
It has long been recognized that one can design Bayesian models that adaptively become more complex when the data demand it. This idea is a core theme in the Bayesian nonparametric literature, although in practice much of the focus has been on clustering and latent feature models (Orbanz and Teh,, 2011; Gershman and Blei,, 2012; Ghahramani et al.,, 2007). A key difference from the settings we consider is that clustering and latent feature problems are evidently complex (with the main question being only how complex), whereas we focus on settings that are plausibly simple, where a parametric model might in fact be sufficient. See Li and Dunson, (2020) for an example of using a baseline nonparametric model as a tool for evaluating parametric models, as we do here.
@@
We briefly describe some other works that have been particularly influential on our approach. Neal, (1995) makes the point that, “There is no provision in the Bayesian framework for changing the model or the prior depending on how much data was collected. If the model and prior are correct for a thousand observations, they are correct for ten observations as well (though the impact of using an incorrect prior might be more serious with fewer observations),” and that when the underlying phenomena is evidently complex the appropriate solution to finding a simpler model outperforming a more complex model is “to design a different complex model that captures whatever aspect of the problem led to the simple model performing well.”
@@
In the context of mixture models, a selling point of infinite mixtures $f(y)=\sum_{k=1}^{\infty}\omega_{k}\,g(y\mid\vartheta_{k})$ is that they can collapse to a single parametric model $g(y\mid\theta)$ when most prior mass is placed on the event $\omega_{1}\approx 1$. For Dirichlet process mixtures (Escobar and West,, 1995), this is accomplished by shrinking the concentration parameter heavily towards zero. For conditional mixture models — such as dependent Dirichlet processes (DDPs, MacEachern,, 2000; Quintana et al.,, 2022) — this allows us to recover standard parametric generalized linear models when appropriate (Shahbaba and Neal,, 2009; Hannah et al.,, 2011).
@@
Our strategy also represents a type of continuous model expansion in the sense of Gelman and Shalizi, (2013), who recommend “forming a larger model that includes both A and B as special cases,” as an alternative to testing Model A against Model B. Like them, we do not formally test the parametric submodel (except as a theoretical device). While they explicitly do not recommend Bayesian nonparametrics as a tool for continuous model expansion, we also do not feel that our proposal is at odds with the additional use of model checking they recommend, and do not view our use of an additional nonparametric component to be a complete solution to model misspecification. Related works include Kamary et al., (2018) and Yao et al., (2018).
1.2 Contributions
@@
We recommend defensive model expansion partly because, under very general conditions, it yields efficient inference when the associated parametric model is correctly specified. The story is particularly simple if we replace the posterior distribution with a fractional posterior (Walker and Hjort,, 2001; O’Hagan,, 1995), which allows us to sidestep the technical testing and entropy conditions required by Ghosal et al., (2000). In fact, the nonparametric component of the prior plays little direct role in our analysis, making the results extremely general.
@@
We begin by presenting general results on posterior concentration for nonparametric models when parametric submodels have sufficient prior mass. With the exception of Theorems 6 and 7, these are straightforward extensions of existing results for fractional posteriors. We prove results for convergence rates, model selection, and asymptotic normality under the assumption that the parametric model is correctly specified. Our model selection results essentially show that (i) we can correctly select the parametric model and (ii) the posterior distribution satisfies a Bernstein–von Mises theorem. We also prove two semiparametric Bernstein–von Mises theorems that apply in settings where model selection consistency is difficult to use, which guarantees efficient posterior concentration of a projection parameter $\theta^{\star}$ around its true value. Such projection parameters are often used to summarize posterior distributions (Woody et al.,, 2021).
@@
We then apply these results to Gaussian process (GP) and Bayesian additive regression tree (BART) models. We show that BART models naturally satisfy the conditions required for a Bernstein–von Mises result (see also Rocková,, 2020), as do “spike-and-GP” priors. While we do not study this here, our semiparametric Bernstein–von Mises results for GPs also have scope for being extended to projection parameters even when the anchoring parametric model is misspecified. These theoretical results are validated on simulated data.
@@
In addition to being theoretically simple, defensive model expansion is easy to implement in practice. For example, because our convergence rate and model selection results apply directly to BART models, existing software for the general BART model of Tan and Roy, (2019) makes it straight-forward to obtain parametric adaptivity with default prior specifications. BART models are especially attractive because, in addition to adapting to parametric structures, they also adjust to locally varying smoothness and the presence of low-order interactions (Linero and Yang,, 2018; Ročková and van der Pas,, 2020; Jeong and Rockova,, 2023).
@@ 1.3 Notation
We assume that $\theta$ decomposes as $(r,\eta)$ where $r\in\mathcal{H}$ is infinite-dimensional and $\eta\in E\subseteq\mathbb{R}^{P}$ is-finite dimensional, such that $\Theta=\mathcal{H}\times E$. The family $\mathcal{F}_{0}=\{F_{\theta}:\eta\in E\ \text{and}\ r=r_{0}\}$ represents a parametric submodel of $\mathcal{F}$. In our examples, $r(\cdot)$ is a mapping $\mathbb{R}^{P}\to\mathbb{R}$ and $r_{0}(\cdot)\equiv 0$. Intuitively, $\eta$ corresponds to the usual parametric component (e.g., regression coefficients), while $r$ is a flexible nonparametric adjustment. We let $\theta_{0}=(r_{0},\eta_{0})$.
@@
Throughout, we mostly assume independent priors on $r$ and $\eta$, i.e., $\Pi=\Pi_{r}\times\Pi_{\eta}$, with $\Pi_{\eta}$ admitting a continuous density $\pi_{\eta}$ on $E$ such that $\pi_{\eta}(\eta_{0})>0$. We also assume that $\mathcal{H}$ is a Hilbert space with inner product $\langle\cdot,\cdot\rangle$, and that there exists a norm $\|\cdot\|_{\mathcal{H}}$ and positive constants $(C(\theta_{0}),\epsilon)$ such that $\max\{K(\theta_{0}\|\theta),V(\theta_{0}\|\theta)\}\leq C(\theta_{0})\{\|\eta-\eta_{0}\|^{2}+\|r-r_{0}\|_{\mathcal{H}}^{2}\}$ for all $\theta\in\Theta$ satisfying $\|\eta-\eta_{0}\|^{2}+\|r-r_{0}\|_{\mathcal{H}}^{2}\leq\epsilon^{2}/C(\theta_{0})$. Here $K(\theta_{0}\|\theta)$ and $V(\theta_{0}\|\theta)$ are Kullback–Leibler type divergences, as defined in Section 3.
@@
Let $Z_{i}=(X_{i},Y_{i})$ with $Y_{i}\mid X_{i}=x]\sim\operatorname{Normal}\{x^{\top}\beta_{0}+r_{0}(x),\sigma_{0}^{2}\},\text{ }X_{i}\sim F_{X},$ where $\theta=(r,\beta,\sigma)$ and $X_{i}\in[0,1]^{D}$ for some $D$. Anchoring to the parametric submodel with $r_{0}(x)\equiv 0$ recovers ordinary linear regression. Our goals are to estimate the regression function $\mu_{0}(x)=x^{\top}\beta_{0}$ and, ideally, recover $\beta_{0}$ efficiently.
@@
We adopt the same framework as Model NR, except $Y_{i}\in\{0,1\}$ and $\mathbb{E}_{\theta}(Y_{i}\mid X_{i}=x)=\mu(x)=(1+\exp\{x^{\top}\beta+r(x)\})^{-1}$ so that $Y_{i}\mid X_{i}=x]$ is a logistic regression and $\theta=(r,\beta)$. Our goal again is to estimate $\mu_{0}(x)$ and estimate $\beta_{0}$ efficiently.
@@ 2.1 General Bayesian Additive Regression Trees Models
For routine use, we recommend the Bayesian Additive Regression Trees (BART, Chipman et al.,, 2010) prior for the nonparametric component $r(\cdot)$. In the context of Model NR the regression function is $\mu(x)=x^{\top}\beta+\sum_{t=1}^{T}g(x;\mathcal{T}_{t},\mathcal{M}_{t})$ where each $g(x;\mathcal{T}_{t},\mathcal{M}_{t})$ is a regression tree, $\mathcal{T}_{t}$ denotes the tree structure, and $\mathcal{M}_{t}=\{\lambda_{\ell t}:\ell=1,\ldots,L_{t}\}$ is the set of leaf node predictions. Tan and Roy, (2019) refer to this as a general BART (GBART) model, and it can be fit using the SoftBart package (Linero,, 2022) using the gsoftbart_regression function. A schematic of the construction of the model is given in Figure 1, which shows a single tree from the ensemble and the associated step function.
Figure 1: Left: an example of a regression tree with input $x=(x_{1},x_{2})$ supported on $[0,1]$. Right: the induced step function on $0,1]^{2}$ .
We now review the prior specification of BART, which is important for our purposes. The default recommendation of Chipman et al., (2010) takes $\lambda_{t\ell}\sim\operatorname{Normal}(0,\sigma^{2}_{\mu}/T)$ with $\sigma_{\mu}=0.5$ after centering and scaling the $Y_{i}$’s to lie between -0.5 and 0.5. The scaling of $\sigma^{2}_{\mu}$ by $T$ is done so that $\operatorname{Var}\{\sum_{t}g(x;\mathcal{T}_{t},\mathcal{M}_{t})\}=\sigma^{2}_{\mu}$, which in turn leads to a Gaussian process limit as $T\to\infty$ (Linero,, 2017). One way to shrink the GBART model to a linear regression is to place a prior on $\sigma_{\mu}$ supported near 0, such as a half-normal or half-Cauchy prior.
Another way for GBART to reduce to a linear regression is for all of the trees in the ensemble to be “empty” in the sense of not splitting on any variables. In Section 3.2 we will show that this suffices for GBART to possess strong theoretical properties when the linear model is correct. Chipman et al., (2010) propose a branching process prior for $\mathcal{T}_{t}$ that (i) decides whether each node at a given depth of the tree will be converted into a branch node with two leaf children; (ii) iterating over the depths; and (iii) terminating if all of the nodes at a given depth are leaf nodes. The probability of a node at depth $d$ being a branch node under the prior is specified to be $q(d)=a/(1+b)^{d}$ where by default $a=0.95$ and $b=2$. In particular, the probability that a tree consists of no splits is (1a)(1-a), and the prior probability that the GBART model reduces to the linear model is therefore $(1-a)^{T}>0$.
@@ 2.2 Gaussian Process Regression
A common approach for anchoring Model NR to a parametric submodel is to use the Gaussian process regression (GP regression) model. GPs are distributions on random functions, and can be used as priors model for regression functions $r(\cdot)$. A GP is determined by a mean function $m(x)$ and a positive semi-definite kernel $\kappa(x,x^{\prime})$, with $\mathbb{E}_{\Pi}r(x)=m(x)$ and $\operatorname{Cov}_{\Pi}\{r(x),r(x^{\prime})\}=\kappa(x,x^{\prime})$. We write $\operatorname{GP}(m,\kappa)$ for the associated Gaussian process. For a review of Gaussian processes and alternative choices of $\kappa(\cdot,\cdot)$, see Rasmussen and Williams, (2006). Importantly, GPs are conjugate priors for $\mu(x)$ in the model $Y_{i}\mid X_{i}=x]\sim\operatorname{Normal}\{\mu(x),\sigma^{2}\}$ when $\kappa(x,x^{\prime})$ is held fixed.
For model NR, we take $r\sim\operatorname{GP}\{0,\sigma^{2}_{\mu}\,\kappa(\cdot,\cdot)\}$. Equivalently, letting $\mu(x)=x^{\top}\beta+r(x)$ we have $\mu\sim\operatorname{GP}(m,\sigma^{2}_{\mu}\,\kappa)$ where $m(x)=x^{\top}\beta$. If $\beta\sim\operatorname{Normal}(0,\sigma^{2}_{\beta})$ this simplifies further to $\mu\sim\operatorname{GP}(0,\sigma^{2}_{\mu}\kappa+\sigma^{2}_{\beta}\kappa_{\text{linear}})$ where $\kappa_{\text{linear}}(x,x^{\prime})=x^{\top}x^{\prime}$. In our illustrations we use the squared exponential function $\kappa_{\text{SE}}(x,x^{\prime})=\exp(-\rho\|x-x^{\prime}\|^{2})$, with the lengthscale $\rho$ controlling the smoothness of samples from the prior.
@@
One option is to place a hyperprior on $\sigma_{\mu}$, with heavy shrinkage towards zero. We call this the scale mixture of GPs approach. Intuitively, this lets the model collapse to the parametric submodel when the nonparametric adjustment is unnecessary by taking $\sigma_{\mu}\approx 0$. When performing inference on projection parameters (see Section 2.4) this strategy is sufficient for most purposes.
@@
A third strategy is to simply use a GP for $r(x)$ and then perform inference on a projection parameter $\eta^{\star}$ obtained from projecting the full model onto the parametric submodel $\mathcal{F}_{0}$. Concretely, we might compute the Kullback-Leibler projection $\theta^{\star}=\arg\min_{\theta\in\mathcal{F}_{0}}K(\theta\|\theta^{\star})$ for each sample $\theta$ from the posterior and, because $\theta_{0}\in\mathcal{F}_{0}$, hopefully $\theta^{\star}$ will concentrate faster around $\theta_{0}$ than $\theta$. This can be used with either of the previous two approaches, but sometimes works on its own as well. When the parametric model is correctly specified, appropriately chosen GP priors can yield efficient inference. However, poorly chosen priors may still lead to inefficient or invalid inference. For example, we will see that even if the posterior converges at a near-parametric rate such as $N^{-1/2}(\log N)^{t}$ for some $t>0$, the bias in estimating $\beta_{0}$ may decay very slowly, making inference inefficient in finite samples. This occurs, for example, when we incorrectly set $\beta=0$ and model $r(x)$ with a squared exponential kernel.

### v2

@@ Abstract
Some applied researchers hesitate to use nonparametric methods, worrying that they will lose power in small samples or overfit the data when simpler models are sufficient. We argue that at least some of these concerns are unfounded when nonparametric models are strongly shrunk toward parametric submodels. We consider expanding a parametric model with a nonparametric component $r(x)$ that is heavily shrunk toward zero. This construction allows the model to adapt automatically: if the parametric model is correct, the nonparametric component disappears, recovering parametric efficiency, while if it is misspecified, the flexible component activates to capture the missing signal. We show that this adaptive behavior follows from simple and general conditions. Specifically, we prove that Bayesian nonparametric models anchored to linear regression, including variants of Gaussian process regression and Bayesian additive regression trees, consistently identify the correct parametric submodel when it holds and give asymptotically efficient inference for regression coefficients. In simulations, we find that the general BART model performs identically to correctly specified linear regression when the parametric model holds, and substantially outperforms it when nonlinear effects are present. This suggests a practical paradigm: defensive model expansion as a safeguard against model misspecification.
@@ 1 Introduction
  • If the parametric model is badly misspecified or omits important signals, expanding into a nonparametric model can improve power. With a flexible “alternative” model where $r(x)\neq 0$, we can also use the estimated size of $r(\cdot)$ as a diagnostic for parametric model misspecification: if $|r(x)|$ is large on average then this is evidence that the parametric submodel fits poorly.
@@ 1.1 Existing Literature
It has long been recognized that one can design Bayesian models that adaptively become more complex when the data demand it. This idea is a core theme in the Bayesian nonparametric literature, although in practice much of the focus has been on clustering and latent feature models (Orbanz and Teh,, 2011; Gershman and Blei,, 2012; Ghahramani et al.,, 2007). A difference from the settings we consider is that clustering and latent feature problems are evidently complex (with the main question being only how complex), whereas we focus on settings that are plausibly simple, where a parametric model might in fact be sufficient. See Li and Dunson, (2020) for an example of using a baseline nonparametric model as a tool for evaluating parametric models, as we do here.
@@
We briefly describe some other works that have been particularly influential on our approach. Neal, (1995) makes the point that, “There is no provision in the Bayesian framework for changing the model or the prior depending on how much data was collected. If the model and prior are correct for a thousand observations, they are correct for ten observations as well (though the impact of using an incorrect prior might be more serious with fewer observations),” and that when the underlying phenomenon is evidently complex the appropriate solution to finding a simpler model outperforming a more complex model is “to design a different complex model that captures whatever aspect of the problem led to the simple model performing well.”
@@
In the context of mixture models, a selling point of infinite mixtures $f(y)=\sum_{k=1}^{\infty}\omega_{k}\,g(y\mid\vartheta_{k})$ is that they can collapse to a single parametric model $g(y\mid\theta)$ when most prior mass is placed on the event $\omega_{1}\approx 1$. For Dirichlet process mixtures (Escobar and West,, 1995), this is accomplished by shrinking the concentration parameter heavily toward zero. For conditional mixture models — such as dependent Dirichlet processes (DDPs, MacEachern,, 2000; Quintana et al.,, 2022) — this allows us to recover standard parametric generalized linear models when appropriate (Shahbaba and Neal,, 2009; Hannah et al.,, 2011).
@@
Our strategy also represents a type of continuous model expansion in the sense of Gelman and Shalizi, (2013), who recommend “forming a larger model that includes both A and B as special cases,” as an alternative to testing Model A against Model B. Like them, we do not recommend formally testing the parametric submodel (except as a theoretical device). While they explicitly do not recommend Bayesian nonparametrics as a tool for continuous model expansion, we also do not feel that our proposal is at odds with the additional use of model checking they recommend, and do not view our use of an additional nonparametric component to be a complete solution to model misspecification. Related works include Kamary et al., (2018) and Yao et al., (2018).
1.2 Outline
@@
We recommend defensive model expansion partly because, under very general conditions, it yields efficient inference when the associated parametric model is correctly specified. The story is particularly simple if we replace the posterior distribution with a fractional posterior (Walker and Hjort,, 2001; O’Hagan,, 1995), which allows us to sidestep the technical testing and entropy conditions required by Ghosal et al., (2000). In fact, the nonparametric component of the prior plays little direct role in our analysis, making the results extremely general. That being said, the main role played by the fractional posterior is to make establishing a posterior contraction rate simpler; the assumption that a fractional posterior is used can often be replaced with directly assuming a posterior concentration rate, for which there are many existing rules for the GP and BART models we consider in this work.
@@
We begin by presenting general results on posterior concentration for nonparametric models when parametric submodels have sufficient prior mass. These are straightforward extensions of existing results for fractional posteriors. We prove results for convergence rates, model selection, and asymptotic normality under the assumption that the parametric model is correctly specified. Our model selection results essentially show that (i) we can correctly select the parametric model and (ii) the posterior distribution satisfies a Bernstein-von Mises theorem.
We also prove two semiparametric Bernstein-von Mises theorems (Theorem 6 and Theorem 7) that apply in settings where model selection consistency is difficult to apply, which guarantee efficient posterior concentration of a projection parameter $\theta^{\star}$ around its true value; when the submodel is correct, this projection parameter will coincide with the parameter of interest. Such projection parameters are often used to summarize posterior distributions (Woody et al.,, 2021). To the best of our knowledge, Theorem 7 is not implied by existing results.
@@
We apply these results to Gaussian process (GP) and Bayesian additive regression tree (BART) models. We show that BART models naturally satisfy the conditions required for a Bernstein-von Mises result (see also Rocková,, 2020), as do “spike-and-GP” priors. While we do not study this here, our semiparametric Bernstein-von Mises results also may be extended to projection parameters even when the anchoring parametric model is misspecified. These theoretical results are validated on simulated data.
@@
In addition to being theoretically simple, defensive model expansion is easy to implement in practice. For example, because our convergence rate and model selection results apply directly to BART models, existing software for the general BART model of Tan and Roy, (2019) makes it straightforward to obtain parametric adaptivity with default prior specifications. BART models are especially attractive because, in addition to adapting to parametric structures, they also adjust to locally varying smoothness and the presence of low-order interactions (Linero and Yang,, 2018; Ročková and van der Pas,, 2020; Jeong and Rockova,, 2023).
@@ 1.3 Notation
We assume that $\theta$ decomposes as $(r,\eta)$ where $r\in\mathcal{H}$ is infinite-dimensional and $\eta\in E\subseteq\mathbb{R}^{P}$ is finite-dimensional, such that $\Theta=\mathcal{H}\times E$. The family $\mathcal{F}_{0}=\{F_{\theta}:\eta\in E\ \text{and}\ r=r_{0}\}$ represents a parametric submodel of $\mathcal{F}$. In our examples, $r(\cdot)$ is a mapping $\mathbb{R}^{D}\to\mathbb{R}$ and $r_{0}(\cdot)\equiv 0$. Intuitively, $\eta$ corresponds to the usual parametric component (e.g., regression coefficients), while $r$ is a flexible nonparametric adjustment. We let $\theta_{0}=(r_{0},\eta_{0})$.
@@
Throughout, we mostly assume independent priors on $r$ and $\eta$, i.e., $\Pi=\Pi_{r}\times\Pi_{\eta}$, with $\Pi_{\eta}$ admitting a continuous density $\pi_{\eta}$ on $E$ such that $\pi_{\eta}(\eta_{0})>0$. We also assume that $\mathcal{H}$ is a Hilbert space with inner product $\langle\cdot,\cdot\rangle$, and that there exists a norm $\|\cdot\|_{\mathcal{H}}$ and positive constants $(C(\theta_{0}),\epsilon)$ such that $\max\{K(\theta_{0}\|\theta),V(\theta_{0}\|\theta)\}\leq C(\theta_{0})\{\|\eta-\eta_{0}\|^{2}+\|r-r_{0}\|_{\mathcal{H}}^{2}\}$ for all $\theta\in\Theta$ satisfying $\|\eta-\eta_{0}\|^{2}+\|r-r_{0}\|_{\mathcal{H}}^{2}\leq\epsilon^{2}/C(\theta_{0})$. Here $K(\theta_{0}\|\theta)$ and $V(\theta_{0}\|\theta)$ are Kullback–Leibler type divergences, as defined in Section 3. Different examples where we apply PPA will use different norms and inner products.
@@
Let $Z_{i}=(X_{i},Y_{i})$ with $Y_{i}\mid X_{i}=x]\sim\operatorname{Normal}\{x^{\top}\beta_{0}+r_{0}(x),\sigma_{0}^{2}\},\text{ }X_{i}\sim F_{X},$ where $\theta=(r,\beta,\sigma)$ and $X_{i}\in[0,1]^{D}$ for some $D$. Anchoring to the parametric submodel with $r_{0}(x)\equiv 0$ recovers ordinary linear regression. Our goals are to estimate the regression function $\mu_{0}(x)=r_{0}(x)+x^{\top}\beta_{0}$ and, ideally, recover $\beta_{0}$ efficiently.
@@
We adopt the same framework as Model NR, except $Y_{i}\in\{0,1\}$ and $\mathbb{E}_{\theta}(Y_{i}\mid X_{i}=x)=\mu(x)=(1+\exp\{-x^{\top}\beta-r(x)\})^{-1}$ so that $Y_{i}\mid X_{i}=x]$ is a logistic regression and $\theta=(r,\beta)$. Our goal again is to estimate $\mu_{0}(x)=\mathbb{E}_{\theta_{0}}(Y_{i}\mid X_{i}=x)$ and estimate $\beta_{0}$ efficiently.
@@ 2.1 General Bayesian Additive Regression Trees Models
For routine use, we recommend the Bayesian Additive Regression Trees (BART, Chipman et al.,, 2010) prior for the nonparametric component $r(\cdot)$. In the context of Model NR, the regression function is $\mu(x)=x^{\top}\beta+\sum_{t=1}^{T}g(x;\mathcal{T}_{t},\mathcal{M}_{t})$ where each $g(x;\mathcal{T}_{t},\mathcal{M}_{t})$ is a regression tree, $\mathcal{T}_{t}$ denotes the tree structure, and $\mathcal{M}_{t}=\{\lambda_{\ell t}:\ell=1,\ldots,L_{t}\}$ is the set of leaf node predictions. Tan and Roy, (2019) refer to this as a general BART (GBART) model, and it can be fit using the SoftBart package (Linero,, 2022) using the gsoftbart_regression function. A schematic of the construction of the model is given in Figure 1, which shows a single tree from the ensemble and the associated step function.
Figure 1: Left: an example of a regression tree with input $x=(x_{1},x_{2})$ supported on $0,1]^{2}$. Right: the induced step function on $0,1]^{2}$ .
We now review the prior specification of BART, which is important for our purposes. The default recommendation of Chipman et al., (2010) takes $\lambda_{t\ell}\sim\operatorname{Normal}(0,\sigma^{2}_{\mu}/T)$ with $\sigma_{\mu}=0.5$ after centering and scaling the $Y_{i}$’s to lie between -0.5 and 0.5. The scaling of $\sigma^{2}_{\mu}$ by $T$ is done so that $\operatorname{Var}\{\sum_{t=1}^{T}g(x;\mathcal{T}_{t},\mathcal{M}_{t})\}=\sigma^{2}_{\mu}$, which in turn leads to a Gaussian process limit as $T\to\infty$ (Linero,, 2017). One way to shrink the GBART model to a linear regression is to place a prior on $\sigma_{\mu}$ supported near 0, such as a half-normal or half-Cauchy prior.
Another way for GBART to reduce to a linear regression is for all of the trees in the ensemble to be “empty” in the sense of not splitting on any variables. In Section 3.2 we will show that this suffices for GBART to possess strong theoretical properties when the linear model is correct. Chipman et al., (2010) propose a branching process prior for $\mathcal{T}_{t}$ that (i) decides whether each node at a given depth of the tree will be converted into a branch node with two leaf children; (ii) iterates over the depths; and (iii) terminates if all of the nodes at a given depth are leaf nodes. The probability of a node at depth $d$ being a branch node under the prior is specified to be $q(d)=a/(1+b)^{d}$ where by default $a=0.95$ and $b=2$. In particular, the probability that a tree consists of no splits is (1a)(1-a), and the prior probability that the GBART model reduces to the linear model (with intercept offset by the constant $r(0)$) is therefore $(1-a)^{T}>0$.
@@ 2.2 Gaussian Process Regression
A common approach for anchoring Model NR to a parametric submodel is to use the Gaussian process regression (GP regression) model. GPs are distributions on random functions, and can be used as priors for regression functions $r(\cdot)$. A GP is determined by a mean function $m(x)$ and a positive semi-definite kernel $\kappa(x,x^{\prime})$, with $\mathbb{E}_{\Pi}r(x)=m(x)$ and $\operatorname{Cov}_{\Pi}\{r(x),r(x^{\prime})\}=\kappa(x,x^{\prime})$. We write $\operatorname{GP}(m,\kappa)$ for the associated Gaussian process. For a review of Gaussian processes and alternative choices of $\kappa(\cdot,\cdot)$, see Rasmussen and Williams, (2006). Importantly, GPs are conjugate priors for $\mu(x)$ in the model $Y_{i}\mid X_{i}=x]\sim\operatorname{Normal}\{\mu(x),\sigma^{2}\}$ when $\kappa(x,x^{\prime})$ is held fixed.
For model NR, we take $r\sim\operatorname{GP}\{0,\sigma^{2}_{\mu}\,\kappa(\cdot,\cdot)\}$. Equivalently, letting $\mu(x)=x^{\top}\beta+r(x)$ we have $\mu\sim\operatorname{GP}(m,\sigma^{2}_{\mu}\,\kappa)$ where $m(x)=x^{\top}\beta$. If $\beta\sim\operatorname{Normal}(0,\sigma^{2}_{\beta}\mathrm{I})$ this simplifies further to $\mu\sim\operatorname{GP}(0,\sigma^{2}_{\mu}\kappa+\sigma^{2}_{\beta}\kappa_{\text{linear}})$ where $\kappa_{\text{linear}}(x,x^{\prime})=x^{\top}x^{\prime}$. In our illustrations we use the squared exponential function $\kappa_{\text{SE}}(x,x^{\prime})=\exp(-\rho\|x-x^{\prime}\|^{2})$, with the lengthscale $\rho$ controlling the smoothness of samples from the prior.
@@
One option is to place a hyperprior on $\sigma_{\mu}$, with heavy shrinkage toward zero. We call this the scale mixture of GPs approach. Intuitively, this lets the model collapse to the parametric submodel when the nonparametric adjustment is unnecessary by taking $\sigma_{\mu}\approx 0$. When performing inference on projection parameters (see Section 2.4) this strategy is sufficient for most purposes.
@@
[…]